OpenStudy (anonymous):

The rectangular prism shown has a volume of 72 m3. This prism is enlarged using a scale factor of 3.

OpenStudy (anonymous):

Is there more to the question ?

OpenStudy (anonymous):

The rectangular prism shown has a volume of 72 m3. This prism is enlarged using a scale factor of 3. What is the volume of the enlarged prism? A. 75 m3 B. 216 m3 C. 648 m3 D. 1944 m3

OpenStudy (anonymous):

\[v=lbh\]

OpenStudy (anonymous):

in this case v=72

OpenStudy (anonymous):

your new volume =(3l)(3b)(3h) =3^3 = 27*72 = ?

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

what's 27 x 72 ?

OpenStudy (anonymous):

1944

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

Right & you're welcome !

OpenStudy (anonymous):

A prism has a surface area of 50 mm2. Which sentence describes the effect of measuring the surface area in square centimeters? A. The surface area stays the same, but the number representing the surface area increases. B. The surface area stays the same, but the number representing the surface area decreases. C. The surface area increases. D. The surface area decreases.